### v1

# TVR: Automotive System Requirement Traceability Validation and Recovery Through Retrieval-Augmented Generation

Abstract.
In automotive software development, as well as other domains, traceability between stakeholder requirements and system requirements is crucial to ensure consistency, correctness, and regulatory compliance. However, erroneous or missing traceability relationships often arise due to improper propagation of requirement changes or human errors in requirement mapping, leading to inconsistencies and increased maintenance costs. Existing approaches do not address traceability between stakeholder and system requirements, rely on open-source data—as opposed to automotive (or any industry) data—and do not address the validation of manual links established by engineers. Additionally, automotive requirements often exhibit variations in the way they are expressed, posing challenges for supervised models requiring training. The recent advancements in large language models (LLMs) provide new opportunities to address these challenges. In this paper, we introduce TVR, a requirement Traceability Validation and Recovery approach primarily targeting automotive systems, leveraging LLMs enhanced with retrieval-augmented generation (RAG). TVR is designed to validate existing traceability links and recover missing ones with high accuracy. We empirically evaluate TVR on automotive requirements, achieving 98.87% accuracy in traceability validation and 85.50% correctness in traceability recovery. Additionally, TVR demonstrates strong robustness, achieving 97.13% in accuracy when handling unseen requirements variations. The results highlight the practical effectiveness of RAG-based LLM approaches in industrial settings, offering a promising solution for improving requirements traceability in complex automotive systems.
copyright: nonecopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; xxx; xxxisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Automotive systems have become increasingly complex, comprising a wide range of integrated components and features such as powertrain systems, chassis systems, infotainment systems, advanced driver assistance systems (ADAS), and electric and autonomous driving systems. These components collaborate through electronic control units (ECUs), sensors, actuators, and software, working seamlessly to deliver performance, safety, and user experience  [^1] https://en.wikipedia.org/wiki/Automotive_electronics. Modern vehicles are thus highly sophisticated multi-system platforms, often referred to as “computers on wheels.”
To ensure high-quality software development, the automotive industry adheres to the Automotive Software Process Improvement and Capability Determination (ASPICE) standard (Idri and Cheikhi, 2016; García-Mireles et al., 2012), which is based on the ISO 3300x norm group (ISO/IEC 33001, 5 03; ISO/IEC 33002, 5 03; ISO/IEC 33003, 5 03; ISO/IEC 33004, 5 03). A fundamental aspect of ASPICE-compliant software development is requirements engineering, where requirements are defined at different levels, including stakeholder requirements and system requirements. System requirements are derived from stakeholder requirements, with traceability links established between them to ensure consistency and correctness, especially in the occurrence of frequent changes.
Maintaining accurate and high-quality traceability links between stakeholder and system requirements is indeed crucial for multiple reasons: 1) Enhancing system consistency; ensuring that requirements are correctly implemented minimizes functional errors and inconsistencies (Wiegers and Beatty, 2013; Pargaonkar, 2023). 2) Preventing error propagation – Detecting incorrect requirement mappings early reduces maintenance costs and avoids large-scale rework (Tufail et al., 2017). 3) Ensuring regulatory compliance, thus complying with ASPICE standards to help improve the safety, reliability, and auditability of automotive software (Qusef et al., 2011).
However, significant challenges have been reported by industry engineers regarding traceability between requirements, for example in the context of Diagnostic Trouble Code (DTC) management, which refer to standardized vehicle error codes indicating malfunctions detected by the diagnostics system (Marscholik and Subke, 2009; Pirasteh et al., 2019). According to system engineers, the primary causes of these challenges are related to resource and time constraints, and include: 1) Improper propagation of requirement changes: When stakeholder requirements are modified (e.g., new requirements are added or existing ones are updated), the corresponding system requirements may not be updated accordingly, leading to outdated or invalid traceability links. If a requirement is deleted or merged, the original traceability link may persist, resulting in incorrect mappings that could cause functional inconsistencies or safety risks. 2) Errors in requirement mapping: Due to human error, system requirements may be incorrectly mapped to unrelated stakeholder requirements, or traceability links are missing.
The goal of this paper is to provide automated, effective means to address the challenges above and thus identify invalid and missing traceability links between DTC requirements. Because requirements are not always expressed consistently within or across projects, we rely on large language models (LLMs), hoping they are robust to such variations.
We are motivated in our choice by the fact that LLMs have demonstrated significant advancements in natural language processing (NLP) (Gao et al., 2024; Karanikolas et al., 2023; Hu and Lu, 2024; Tang et al., 2023) and software engineering tasks (Gao et al., 2024; Hu and Lu, 2024; Karanikolas et al., 2023), especially code generation (Du et al., 2024; Mu et al., 2024; Gao et al., 2023a; Chen et al., 2024; Zhang et al., 2024; Tang et al., 2023), and code summarization (Nam et al., 2024; Ahmed et al., 2024; Geng et al., 2024). Further, LLMs have also shown to be effective for traceability link recovery (Fuchß et al., 2025; Rodriguez et al., 2023).
To address traceability issues and meet the practical needs of our industrial partner, we propose a Retrieval-Augmented Generation (RAG) enhanced approach: Traceability Validation and Recovery (TVR). We investigate and leverage many state-of-the-art (SOTA) LLMs—with Claude Sonnet 3.5 (Anthropic, 2024) yielding the best results—to verify whether a traceability link is valid and recover missing links between requirements. We evaluated TVR on 2,312 DTC requirement pairs from an automotive system, achieving an overall accuracy of 98.87%. To investigate its robustness against variations in the way requirements are written (as detailed in Section 2), we evaluated TVR on new requirement variations. The results show that TVR still maintains 97.13% in accuracy, demonstrating strong robustness. Additionally, to address the challenge of missing links, we applied TVR to traceability link recovery, successfully identifying 502 missing traceability links with an accuracy of 85.50%.
Our contributions include:
  • Proposing TVR, a RAG-enhanced LLM approach for validating traceability links between automotive requirements.
  • Evaluating the robustness of TVR on unseen requirement variations.
  • Applying TVR to recover traceability links.
  • Investigating 13 SOTA LLMs across five prompt engineering strategies, as well as the impact of different similarity measures and the number of examples on RAG performance.
The remainder of this paper is structured as follows: Section 2 defines the industrial problems we address and provides the context. Section 3 introduces our TVR approach. The study design is detailed in Section 4, followed by an analysis of the experimental results in Section 5, and a discussion in Section 6. Section 7 examines threats to validity. Section 8 reviews the state of the art in requirements traceability research and contrasts it with our contributions. Finally, we conclude the paper in Section 9.
2. Problem Definition
This work aims to support traceability validation and recovery between stakeholder requirements and system requirements for DTCs within automotive systems. This section provides an overview of DTCs and requirements traceability, with a particular focus on stakeholder and system requirements. Though our terminology and sanitized examples come from the automotive domain, many other critical domains, which typically need to comply with functional safety standards, have similar challenges and concepts.
2.1. Diagnostic Trouble Code Requirements
2.1.1. Diagnostic Trouble Code
Diagnostic trouble codes (DTCs) (also known as fault codes) are standardized codes used in automotive systems to identify and diagnose issues within a vehicle’s electronic control units (ECUs) (Marscholik and Subke, 2009; Pirasteh et al., 2019). These codes are generated when a malfunction is detected by the On-Board Diagnostics (OBD) system, which monitors various components such as the engine, transmission, and emission systems. Example conditions that trigger a DTC are implausible or erroneous signal values or signals that were not received (Theissler, 2017). DTCs typically consist of a letter (indicating the system, e.g., P for Powertrain) followed by four digits that provide specific details about the fault. Technicians and diagnostic tools use these codes to pinpoint problems efficiently, aiding in vehicle repair and maintenance.
In modern automotive systems, DTC requirements are expressed at various levels of abstraction: stakeholder requirements and system requirements. These requirements define the conditions under which faults are detected, recorded, and communicated within ECUs, ensuring that the system operates effectively and meets both technical and regulatory standards. In this study, we focus on the requirements pertaining to the setting and clearing of DTCs, corresponding to the conditions for so-called mature and demature DTCs, respectively, which are further described in Section 2.1.3.
2.1.2. DTC Stakeholder Requirements
Stakeholder requirements are the high-level needs and expectations of all parties involved or affected by the automotive system. As shown in Figure 1, DTC stakeholder requirements include the following basic elements:
  • Trigger Condition: Specify the conditions under which a DTC is set or cleared.
  • Input Message: Define the message that triggers the setting or clearing of the DTC.
  • Mitigation Action: Include specific actions for setting the DTC to “Present” or “Not Present”.
  • Validation Rules: Establish validation rules to ensure that DTC setting and clearing operations comply with design standards.
  • Reference Documents: Provide relevant standards and documents supporting the setting and clearing of the DTC.
[…]
Figure 1. Fictitious Variations of Stakeholder Requirements
[…]
When put together, all these basic elements constitute a complete DTC stakeholder requirement. In practice, a stakeholder requirement may encompass multiple independent requirements as it may include both setting and clearing DTC requirements, which follow the same template. In such cases, we can decompose it into atomic requirements based on these fundamental elements, ensuring that each atomic requirement is evaluated independently for traceability.
2.1.3. DTC System Requirements
System requirements define the detailed specifications that the automotive system must meet to satisfy stakeholder requirements. These requirements translate stakeholder expectations into specific, actionable objectives for the system.
In our context, the elements of a DTC system requirement include Name, Number, Description, Priority, Enable Condition, Mature, Mature Time, Demature, Demature Time, and others. Among these, the most critical elements are Mature and Demature. Mature specifies the conditions that must be met for a DTC to be set, while Demature defines the conditions required for the DTC to be cleared. Mature and Demature should be negations of each other.
Figure 2 presents an example of a Mature condition from a system requirement. Both Mature and Demature processes start by verifying that the relevant components are enabled. This ensures that the system is actively monitoring the specified components and is prepared to conduct further checks if the components are available and operational. Next, the system verifies proper communication and ensures that expected data has been received, confirming that the system is functioning correctly. If a malfunction is detected, the system sets the corresponding DTCs; otherwise, the DTCs are cleared.
[…]
Figure 2. Example of Sanitized DTC System Requirement
[…]
2.1.4. DTC Types
There are different types of ECU failures, corresponding to various DTC types, and can be broadly categorized into: internal failures and network failures (Palai, 2013; Theissler, 2017). Network failures can be further divided into two types (Palai, 2013; Theissler, 2017):
  • Lost Communication: This occurs when an ECU fails to receive an expected message from a source ECU. The absence of an expected message is categorized as a Missing Message failure mode.
  • Implausible Data: This occurs when an ECU receives an expected message but detects untrustworthy data that is inconsistent, unrealistic, or beyond the expected domain.
Different types of DTCs follow distinct patterns. In this study, we primarily focus on the two types of network failures due to data availability: “Lost Communication” and “Implausible Data”. Figure 1 illustrates, with sanitized examples, the variations we observed across stakeholder requirements for network failures. Variation 1 represents a “Lost Communication” failure, where the expected message (“MESSAGE_1”) was missing for a certain number of cycles. Variations 2, 3, and 4 correspond to different cases of “Implausible Data” faults, where the ECU receives untrustworthy data. In addition, Variations 1, 2, and 3 describe conditions for “Mature”, which involve setting DTCs, while Variation 4 describes “Demature”, which involves clearing DTCs.
2.2. Requirements Traceability
Requirements traceability refers to the ability to track and document the lifecycle of a requirement—both forward and backward—from its origin through development, implementation, and usage (Gotel and Finkelstein, 1994). In the context of automotive systems, traceability helps ensure that stakeholder needs are effectively aligned with the system design. It plays a critical role in confirming that the system meets customer expectations and regulatory standards. Moreover, traceability provides a clear audit trail for quality assurance, helping to detect gaps, inconsistencies, or changes in requirements throughout the development process.
As discussed in Section 2.1, DTC requirements come in different types, each following distinct patterns. Moreover, both stakeholder and system requirements exhibit variations in the way they are expressed, with the possibility of encountering unseen variations in the future. This presents a significant challenge: A general solution capable of handling diverse and unseen variations is essential. Traditional NLP approaches (Rahimi and Cleland-Huang, 2018; Tufail et al., 2017; Charalampidou et al., 2021), based on supervised learning of observed variations, are challenged by unseen variations. Moreover, a comparison of stakeholder requirements (Figure 1) and system requirements (Figure 2) raises another challenge: Our traceability validation problem cannot be simply achieved by calculating text similarity or finding word overlaps. Since the system requirement involves checking whether multiple messages or signals are missing, while a stakeholder requirement only addresses the handling of a single message, a system requirement may be traced to several stakeholder requirements, each one covering a message or signal in the system requirement. Further, as shown in Figure 1 and Figure 2, messages and signals only differ slightly in terms of character strings.
To address these two challenges, this paper proposes TVR, a method for verifying requirements traceability for automotive systems. TVR leverages the exceptional natural language understanding capabilities of LLMs. Instead of relying on the general concept of traceability links (Guo et al., 2024), we ask LLMs to verify whether our system requirement covers the message or signal specified in the stakeholder requirement. We guide the LLM to focus specifically on the message or signal, rather than other parts of the requirement, to avoid confusion.
3. TVR
TVR is a retrieval-augmented approach that leverages LLMs for requirements traceability validation and recovery. Given that observed variations are specific to the domain and context, we do not retrieve relevant knowledge from an external knowledge base. Instead, inspired by (Li et al., 2023; Wu et al., 2024), we retrieve similar requirement pairs to serve as demonstrations in the prompt. Given a language model $G$, a training set
$D_{\text{train}}=\left\{\begin{aligned} &\langle stakeReq_{i},sysReq_{i},label_{i}\rangle\mid i=1,2,\dots,N,\\ &\text{where }label_{i}\in\{Yes,No\}\end{aligned}\right\}$
and a test case
$x_{\text{test}}=\{stakeReq_{\text{test}},sysReq_{\text{test}}\},$
TVR retrieves similar examples from $D_{\text{train}}$ to assist $G$ in predicting the label for $x_{\text{test}}$.
The overall workflow of TVR, illustrated in Figure 3, consists of two key components: a retriever and a generator. The retriever identifies the most similar examples from $D_{\text{train}}$ based on their similarity to $x_{\text{test}}$ and provides them as input to the generator. The generator then utilizes these retrieved examples to assess the validity of the traceability link between $stake_{\text{test}}$ and $sysReq_{\text{test}}$.
Figure 3. Overall Framework of TVR.
[…]
3.1. Retriever
[…]
3.2. Generator
The generator plays a crucial role in evaluating the traceability between a given pair $\langle stakeReq_{\text{test}},sysReq_{\text{test}}\rangle$. This process leverages retrieval-augmented, in-context learning powered by LLMs, e.g., Claude 3.5 Sonnet  [^4] https://aws.amazon.com/bedrock/claude/, the best model as presented in Section 5. It makes use of the top $k\times 2$ similar examples obtained by the retriever. These retrieved examples are incorporated into the generator’s prompt as contextual examples to guide its response generation. Using this augmented prompt, the generator analyzes the relationship between the stakeholder requirement ($stakeReq_{\text{test}}$) and the system requirement ($sysReq_{\text{test}}$) and generates a response indicating whether their traceability link is valid or not.
As shown in Figure 4, the input of the generator includes:
  1. (1)
    Instruction: The specific task that the model needs to perform.
    Rationale: We define our task as verifying whether a system requirement covers a stakeholder requirement. While traceability validation is a broad concept, we narrow its scope in our context to focus specifically on whether the system requirement covers the message or signal in the stakeholder requirement. This refinement is necessary because a single system requirement may correspond to multiple stakeholder requirements in our case.
    Our Prompt: See Figure 4, Lines 1–2.
  2. (2)
    Context: External information or additional context that can steer the model to better responses.
    Rationale: In our context, we provide the $k\times 2$ most similar examples, including $k$ valid and $k$ invalid ones, to help LLMs learn to understand them.
    Our Prompt: See Figure 4, Line 3.
  3. (3)
    Input Data: the input or question we aim to answer.
    Rationale: We use XML tags to clearly separate different parts of the input data and ensure the prompt is well structured  [^5] https://docs.anthropic.com/en/docs/build-with-claude/prompt-engineering/use-xml-tags.
    Our Prompt: See Figure 4, Lines 4–7.
  4. (4)
    Output Indicator: The type or format of the output.
    Rationale: For experimental purposes, we only require the model to output a predicted label, so the model only needs to respond with “Yes” or “No.” Note that in practice, if an explanation from LLMs is required, this sentence should be adjusted to enable LLMs to generate a step-by-step reasoning and analysis process.
    Our Prompt: “only respond with either ‘Yes’ or ‘No’.” (Line 4)
With the above prompt, the generator analyzes the input requirement pair according to the instructions and provided contextual examples, and then responds with either “Yes” or “No” for the given pair $\langle stakeReq_{\text{test}},sysReq_{\text{test}}\rangle$.
[…]
Figure 4. Input Prompt.
3.3. Traceability Recovery
To recover potentially missing traceability links between any pair of stakeholder requirement and system requirement in the dataset, we first pair every stakeReq with every sysReq that lacks a trace link. This results in a number of pairs approaching the Cartesian product of the two sets (excluding existing traceability links), which would be computationally expensive and time-consuming to analyze.
To address this challenge, we apply the following preprocessing steps to reduce computational overhead:
  1. Step 1:
    Variations Matching: As discussed in Section 2, in the studied dataset, stakeholder requirements can be grouped into four distinct variations, each following a specific template, while system requirements are classified into two categories characterized by unique syntactic patterns. A traceability link can only exist between requirements that belong to the same type of DTC. For example, a stakeholder requirement for “Lost Communication” can only be linked to a system requirement of the same DTC type. By excluding cross-category mismatches, we substantially reduce the number of pairs being considered.
  2. Step 2:
    Condition Matching: Each atomic stakeholder requirement corresponds exclusively to either a mature condition or a demature condition in the system requirement. Based on this distinction, we group stakeReq and sysReq accordingly and only match those within the same condition type, thus further eliminating irrelevant pairs.
  3. Step 3:
    Message Overlap Matching: If a stakeholder requirement and a system requirement do not share any message overlap, there is no traceability between them. To check this, we first tokenize the stakeholder and system requirements, remove stop words (customized for our domain), and extract the messages. We then compare these messages and retain only the pairs where at least one common message exists between them. This step significantly reduces the number of pairs requiring further validation.
Through these three steps, we effectively reduce the number of candidate pairs while preserving all potential missing links. The filtered pairs are then used to validate the following hypothesis (H):
[…]
We use TVR to validate this hypothesis. If H is confirmed for a pair, it implies that the traceLink is missing and should be added to the dataset. Conversely, if H is not confirmed, it indicates that no traceability link exists between stakeReq and sysReq. Through this process, we systematically recover missing traceability links in the dataset.
4. Study Design
In this section, we present our research questions, the LLMs and prompt engineering strategies evaluated, the dataset used for validating and recovering requirements traceability, and the evaluation metrics employed to assess the performance of TVR.
4.1. Research Questions
  1. RQ1:
    What is the performance of SOTA LLMs, with zero-shot prompting, on requirements traceability validation?
    This RQ aims to evaluate and compare the performance of available SOTA LLMs, including Llama, Claude, Titan, and Mistral, with zero-shot prompting.
  2. RQ2:
    What is the best prompt strategy using the best LLM for requirements traceability validation?
    This RQ investigates and compares the performance of various prompt engineering strategies applied to requirements traceability validation, including zero-shot, few-shot, chain-of-thought (CoT), self-consistency, and RAG.
  3. RQ3:
    How robust is TVR for requirements traceability validation in the presence of unseen requirements variations?
    One of our main motivations in relying on LLMs is that we expect them to be more robust to such variations. Due to the numerous variations that are typically present in stakeholder requirements, this RQ primarily evaluates the performance of TVR across unseen variations in stakeholder requirements.
  4. RQ4:
    What is the performance of TVR on traceability link recovery between stakeholder requirements and system requirements?
    LLMs can not only be used for traceability validation but also for missing links recovery by determining if there is valid traceability between any two requirements that are not linked. This research question leverages our TVR approach to recover missing links and evaluates its accuracy.
4.2. Large Language Models and Prompts Engineering
4.2.1. LLMs
In this study, due to our industry partner’s data privacy concerns, we were unable to use LLMs such as the GPT series (OpenAI, 2023a, b). Instead, we were provided with 13 SOTA LLMs available through the Amazon Bedrock service, which include Claude, Llama, Mistral, and Titan. Extensive experiments were conducted using these LLMs.
  • Claude (https://claude.ai/, 2023) is a family of LLMs developed by Anthropic. These models offer several notable advantages: (1) a larger context window, (2) strong performance across various tasks, and (3) not storing any user input or output data.  [^6] https://www.ibm.com/think/topics/claude-ai The evaluated models include Claude 3.5 Sonnet, Claude 3 Sonnet, Claude 2, Claude Instant, and Claude 3 Haiku.
  • Llama (https://www.llama.com/, 2023) is a family of autoregressive LLMs released by Meta AI, which provides models ideal for natural language tasks like Q&A and reading comprehension. The evaluated models are Llama 3 8B and Llama 3 70B.
  • Mistral AI (https://mistral.ai/, 2023) provides models with publicly available weights supporting a variety of use cases from text summarization, text classification, and text completion to code generation and code completion. The evaluated models include Mistral 7B, Mixtral 8x7B, and Mistral Large 2402.
  • Amazon Titan (https://aws.amazon.com/bedrock/amazon models/titan/, 2023) foundation models are a family of foundation models pre-trained by AWS on large datasets, making them powerful, general-purpose models built to support a variety of use cases. The evaluated models include Titan Text Premier, Titan Text Express, and Titan Text Lite.
4.2.2. Prompt Engineering
We select the following SOTA prompt engineering strategies for evaluating LLMs:
  • Zero-shot Prompting (Brown et al., 2020). Zero-shot prompting directly instructs the model to complete a task without providing demonstrations or additional context.
  • Chain-of-Thought Prompting (CoT) (Wei et al., 2022). CoT enhances the complex reasoning capabilities of LLMs by incorporating intermediate reasoning steps.
  • Few-shot Prompting (Brown et al., 2020). Few-shot prompting improves model performance by including a small number of task demonstrations in the prompt. These demonstrations serve as conditioning, guiding the model’s response for subsequent inputs (Brown et al., 2020; Work, [n. d.]).
  • Self-Consistency (Wang et al., 2022). Self-consistency refines CoT prompting by sampling multiple diverse reasoning paths and selecting the most consistent answer. Self-consistency improves the reliability of CoT prompting, especially for arithmetic and commonsense reasoning, by leveraging agreement across multiple reasoning trajectories (Wang et al., 2022).
  • Retrieval-Augmented Generation (RAG) (Lewis et al., 2020). RAG enhances LLMs’ capability of solving knowledge-intensive tasks by integrating information retrieval with text generation. It retrieves relevant external knowledge sources to provide additional context, enabling models to generate more accurate and up-to-date responses without requiring full retraining (Gao et al., 2023b; Riedel et al., 2020).
Note that in our experiments, each prompting strategy builds upon the previous one. The CoT reasoning extends Zero-Shot, while few-shot prompting enhances CoT by incorporating examples in the prompt. The self-consistency strategy further refines few-Shot prompting by applying a majority voting method. RAG improves few-shot prompting by selecting more relevant examples.
4.3. Dataset
4.3.1. Data Collection
To evaluate the performance of LLMs in autonomous requirements traceability validation and recovery, we collected a dataset of DTC requirements from our industry partner. We began by gathering high-quality DTC system requirements, including the “Lost Communication” and “Implausible Data” DTC system requirements, which are among the most common and critical types of DTC requirements. Subsequently, we identified the stakeholder requirements linked to these system requirements, yielding a total of 1,320 stakeholder requirements.
4.3.2. Data Cleaning
After obtaining the system requirements and stakeholder requirements, we performed data cleaning to ensure the dataset’s high quality.
First, if a stakeholder requirement contained multiple atomic requirements, we divided it (based on Section 2) so that each stakeholder requirement contained only a single atomic requirement. We then established the following two constraints for dataset cleaning:
  1. (1)
    All stakeholder requirements must be complete.
  2. (2)
    Every DTC system requirement must have been linked to at least one stakeholder requirement.
To check the completeness of stakeholder requirements, we defined regular expressions to validate their structure 777while we acknowledge that more advanced approaches for completeness checking exist, this is not the primary focus of our study. Specifically, each atomic requirement must include a trigger condition, an input message, and a mitigation action. We filtered out both stakeholder and system requirements that did not meet the above two constraints. In practice, such requirements should be flagged as issues requiring correction and must be addressed before performing requirement traceability validation.
Finally, the dataset includes a total of 1,320 atomic stakeholder requirements linked to 48 system requirements through 2,132 existing traceability links. The dataset contains 24 “Lost Communication” system requirements and 24 “Implausible Data” system requirements.
4.3.3. Data Annotation
We manually annotated 2,132 traceability links between stakeholder requirements and system requirements pairs following the guidelines provided by system engineers employed by our industry partner. 1,913 traceability links were confirmed as valid, while 219 were identified as invalid, indicating that no traceability link should exist between them.
4.4. Evaluation Metrics
4.4.1. Traceability Validation
To assess the performance of TVR in autonomous requirements traceability validation, we employ standard metrics, including accuracy, precision, recall, and F1 score. $TP$ is the number of traceability links that are correctly identified as valid. $FN$ is the number of traceability links that are incorrectly identified as invalid. $FP$ is the number of traceability links that are incorrectly identified as valid. $TN$ is the number of traceability links that are correctly identified as invalid.
Accuracy is the proportion of correctly identified traceability links (both valid and invalid) out of all traceability links. Precision is the proportion of correctly identified valid links ($TP$) out of all links identified as valid by the model ($TP$ + $FP$). It measures how reliable the model’s predictions of valid links are. Recall is the proportion of correctly identified valid links ($TP$) out of all actual valid links in the ground truth ($TP$ + $FN$). It measures the model’s ability to find all the valid links. F1 score is the harmonic mean of precision and recall. It balances the trade-off between precision and recall, providing a single measure of the model’s performance.
4.4.2. Traceability Recovery
To evaluate the performance of TVR in autonomous requirements traceability recovery, we manually verify their correctness (Formula 1). The evaluation metric is the proportion of correct traceability links, calculated as the number of links confirmed as correct through human verification divided by the total number of links identified by the model.
(1) $Correctness=\frac{Number\,of\,Correct\,Links\,Verified\,by\,Humans}{Total\,Number\,of\,Links\,Predicted\,by\,the\,Model}$
5. Experiment Results
5.1. Performance of SOTA LLMs on Requirements Traceability Validation (RQ1)
Approach. This RQ examines the performance of TVR using SOTA LLMs with zero-shot prompting for requirements traceability validation, focusing on models available to our industry partner, including Claude, Llama, Mistral, and Amazon Titan. Specifically, the zero-shot prompting strategy involves providing the task description in the prompt without giving the model any examples.
Results. The experimental results are summarized in Table 1. Claude models generally outperform other model series, with Claude 2 (90.57%), Claude Instant (84.43%), and Claude 3.5 Sonnet (79.55%) achieving the highest accuracy scores. In contrast, Llama 3 70B (10.19%) and Mistral 8×7B (17.23%) show notably lower accuracy, indicating weaker generalization capabilities for this task.
For valid pairs, most models exhibit high precision (above 90%), but recall varies significantly: Mistral Large 2402 achieves the highest precision (100%) but has a recall of only 0.73%, leading to a low F1 score of 1.44%. This suggests the model is highly conservative, prioritizing precision over recall, but missing many valid pairs. Claude 2 achieves the best F1 score (95.04%), with well-balanced precision (90.64%) and recall (99.90%), demonstrating strong recall and overall predictive stability. Llama 3 70B and Mistral Large 2402 exhibit poor recall (only 0.95% and 0.73%, respectively), indicating difficulty in capturing valid cases. Their F1 scores (1.88% and 1.44%) further confirm their limitations. Within the Claude series, Claude 2 outperforms other models, achieving the highest F1 score (95.04%), while Claude Instant follows with an F1 score of 91.51%.
However, detecting invalid pairs proves significantly more challenging, as evidenced by the lower F1 scores across all models. Claude 2 has the highest invalid precision (71.43%) but suffers from extremely low recall (2.45%), resulting in an F1 score of only 4.74%, indicating it misses most invalid pairs. Mistral Large 2402 achieves the highest recall for invalid cases (100.00%) but at a cost of low precision (9.63%), leading to an F1 score of 17.57%. This suggests it identifies all invalid pairs but at the expense of many false positives. Among weaker performers, Claude Instant has an invalid recall of only 4.90% and an F1 score of 5.68%. Llama 3 70B achieves high recall (96.53%) but with low precision (9.44%), leading to a low F1 score of 17.20%.
The reasons for the models’ poor performance include: (1) Even when explicitly instructed to focus on determining whether a message or signal from the stakeholder requirement is covered by the linked system requirement, without domain knowledge, LLMs still struggle to accurately understand and identify the key relevant message or signal in the requirements but rather focus on the wrong aspects, leading to incorrect predictions. (2) LLMs tend to check for content consistency between requirements and often misjudge the two requirement descriptions as inconsistent due to differences in wording or structure. Since invalid pairs are few in number, the small denominator in score calculations, such as precision and recall, can lead to unreliable scores.
Answering RQ1: Results suggest unsatisfactory accuracy for all models with zero-shot prompting for requirements traceability validation, especially for invalid links, thus underscoring the necessity of employing more advanced prompt engineering strategies.
Table 1. Experimental Results of SOTA Models with Different Prompting Strategies.
Prompt Model Acc(%) Valid Invalid
Pre(%) Recall(%) F(%) Pre(%) Recall(%) F(%)
Zero-shot Claude 3.5 Sonnet 79.55 93.42 83.25 88.04 21.98 44.61 29.45
Claude 3 Sonnet 32.08 96.51 25.83 40.75 11.51 91.18 20.44
Claude 2 90.57 90.64 99.90 95.04 71.43 2.45 4.74
Claude Instant 84.43 90.22 92.84 91.51 6.76 4.90 5.68
Claude 3 Haiku 49.72 91.00 49.27 63.93 10.11 53.92 17.03
Llama 3 8B 72.70 91.09 77.39 83.68 11.74 28.43 16.62
Llama 3 70B 10.19 72.00 0.95 1.88 9.44 96.53 17.20
Mistral 7B 57.09 91.96 57.45 70.72 12.08 53.77 19.72
Mixtral 8x7B 17.23 89.76 9.62 17.38 9.43 89.55 17.07
Mistral Large 2402 10.23 100.00 0.73 1.44 9.63 100.00 17.57
Titan Text Premier 20.68 92.17 13.43 23.45 9.83 89.22 17.71
Titan Text Express 49.81 90.17 49.95 64.29 9.30 48.53 15.62
Titan Text Lite 70.83 91.49 74.69 82.24 12.54 34.31 18.37
CoT Claude 3.5 Sonnet 76.31 98.70 $\uparrow$ 74.79 85.10 27.57 $\uparrow$ 90.69 $\uparrow$ 42.29 $\uparrow$
Claude 3 Sonnet 22.92 95.24 15.60 26.81 10.33 92.57 $\uparrow$ 18.59
Claude 2 66.47 94.39 $\uparrow$ 66.79 78.22 17.24 63.54 $\uparrow$ 27.12 $\uparrow$
Claude Instant 86.16 $\uparrow$ 90.36 $\uparrow$ 94.81 $\uparrow$ 92.53 $\uparrow$ 8.26 $\uparrow$ 4.41 5.75 $\uparrow$
Claude 3 Haiku 45.55 $\uparrow$ 94.97 $\uparrow$ 41.62 57.88 13.44 $\uparrow$ 80.42 $\uparrow$ 23.03 $\uparrow$
Llama 3 8B 86.25 $\uparrow$ 90.56 94.73 $\uparrow$ 92.59 $\uparrow$ 6.00 3.30 4.26
Llama 3 70B 35.32 $\uparrow$ 98.17 $\uparrow$ 28.94 $\uparrow$ 44.70 $\uparrow$ 12.51 $\uparrow$ 94.97 22.11 $\uparrow$
Mistral 7B 73.15 $\uparrow$ 92.35 $\uparrow$ 76.63 $\uparrow$ 83.76 $\uparrow$ 15.65 $\uparrow$ 40.59 22.59 $\uparrow$
Mixtral 8x7B 14.53 96.43 $\uparrow$ 5.80 10.94 9.82 $\uparrow$ 97.95 $\uparrow$ 17.84 $\uparrow$
Mistral Large 2402 10.27 $\uparrow$ 100.00 0.78 $\uparrow$ 1.54 $\uparrow$ 9.64 $\uparrow$ 100.00 17.58 $\uparrow$
Titan Text Premier 25.70 $\uparrow$ 92.57 $\uparrow$ 19.40 $\uparrow$ 32.08 $\uparrow$ 10.07 $\uparrow$ 85.29 18.01 $\uparrow$
Titan Text Express 54.41 $\uparrow$ 90.78 $\uparrow$ 55.19 $\uparrow$ 68.65 $\uparrow$ 10.00 $\uparrow$ 47.06 16.49 $\uparrow$
Titan Text Lite 69.93 $\uparrow$ 90.70 $\uparrow$ 74.38 81.73 10.34 27.94 15.10
Few-shot Claude 3.5 Sonnet 97.47 $\uparrow$ 98.70 98.50 $\uparrow$ 98.60 $\uparrow$ 86.06 $\uparrow$ 87.75 86.89 $\uparrow$
Claude 3 Sonnet 90.81 $\uparrow$ 94.00 95.95 $\uparrow$ 94.97 $\uparrow$ 52.44 $\uparrow$ 42.16 46.74 $\uparrow$
Claude 2 41.44 97.88 $\uparrow$ 36.01 52.66 13.29 92.65 $\uparrow$ 23.25
Claude Instant 90.95 $\uparrow$ 96.82 $\uparrow$ 93.05 94.90 $\uparrow$ 51.97 $\uparrow$ 71.08 $\uparrow$ 60.04 $\uparrow$
Claude 3 Haiku 92.73 $\uparrow$ 93.65 98.65 $\uparrow$ 96.08 $\uparrow$ 74.26 $\uparrow$ 36.76 49.18 $\uparrow$
Mistral 7B 90.01 $\uparrow$ 90.62 99.22 $\uparrow$ 94.73 $\uparrow$ 28.57 $\uparrow$ 2.94 5.33
Mixtral 8x7B 30.66 $\uparrow$ 99.13 $\uparrow$ 23.56 $\uparrow$ 38.07 $\uparrow$ 11.90 $\uparrow$ 98.03 $\uparrow$ 21.23 $\uparrow$
Mistral Large 2402 61.35 $\uparrow$ 98.94 57.88 $\uparrow$ 73.04 $\uparrow$ 19.12 $\uparrow$ 94.12 31.79 $\uparrow$
Self-Consistency Claude 3.5 Sonnet 97.61 $\uparrow$ 98.4 98.96 $\uparrow$ 98.68 $\uparrow$ 89.64 $\uparrow$ 84.8 87.15 $\uparrow$
TVR (RAG) Claude 3.5 Sonnet 98.87 $\uparrow$ 99.28 $\uparrow$ 99.48 $\uparrow$ 99.38 $\uparrow$ 95.00 $\uparrow$ 93.14 $\uparrow$ 94.06 $\uparrow$
5.2. Best Prompt Strategy for Requirements Traceability Validation (RQ2)
Approach. This RQ explores the performance of different prompt engineering strategies for the requirement traceability validation task, specifically including CoT, few-shot, self-consistency, and RAG.
For few-shot prompting, we consider all four distinct variations of stakeholder requirements presented in our context, each corresponding to either mature or demature conditions of the linked system requirements. Additionally, the traceability link between each stakeholder requirement and system requirement pair can be either valid or invalid. Given these factors, our dataset contains 16 possible combinations. To maximize the effectiveness of LLMs, we randomly selected one example from each combination, resulting in a total of 16 examples for the prompt. However, the available Llama series models (Llama 3 8B and Llama 3 70B) and Amazon Titan series models (Text Premier, Text Express, and Text Lite) have a limited token capacity (8K), which is insufficient for a 16-shot prompt. Consequently, we excluded these five models from our few-shot prompt experiments, leaving eight models for requirements traceability validation. For these models, we include all 16 examples in the prompt.
Based on the results shown in Table 1, we select the best-performing model (i.e., Claude 3.5 Sonnet) for few-shot prompting and then evaluate self-consistency on this model. Specifically, we run the model ten times and apply a majority voting method, selecting the most frequent output among the ten results as the final output of the model (Wang et al., 2022).
For RAG, we also use Claude 3.5 Sonnet and then use the approach described in Section 3 to retrieve the $k\times 2$ most similar examples as part of the prompt. Due to the limited amount of manually labeled ground truth data, we employ leave-one-out cross-validation (Hastie et al., 2009) in our experiment to make efficient use of the data and ensure realistic estimates. In each iteration, one requirements pair is used as test, while the remaining $n-1$ pairs serve as the retrieval database. We acknowledge that, in practice, companies can obtain a larger volume of high-quality labeled data to enhance the retriever’s database. For the retriever component, we experimentally compare two similarity measures, namely cosine similarity and Euclidean distance, and evaluate the impact of different values of k, ranging from 1 to 8.
Results. The experiment results of different models with various prompt engineering strategies are shown in Table 1. We use $\uparrow$” to highlight improved scores compared to the previous prompt strategy in the table.
The experimental results indicate that incorporating CoT reasoning significantly enhances recall for both valid and invalid cases across most models. Claude 3.5 Sonnet, Claude 2, and Claude 3 Haiku exhibit substantial improvements in invalid recall, leading to higher F1 scores. Meanwhile, Llama 3 8B, Llama 3 70B, Mistral 7B, and Amazon Titan Text Premier show notable gains in valid F1 scores. However, for Claude 3 Sonnet, Llama 3 8B, and Amazon Titan Text Lite, invalid F1 scores decreased due to a drop in precision when detecting invalid pairs.
Comparing the results of few-shot prompting with CoT prompting, we can observe that adding examples to the prompt effectively improves the F1 score for most models in both categories (valid and invalid pairs), except for Claude 2 and Mistral 7B. Apart from Claude 2, all models show a significant improvement in precision for the invalid category. This demonstrates that including examples in the prompt helps models better distinguish the valid and invalid requirement pairs, with significant improvements in the precision for the invalid pairs. However, the drawback of adding too many examples is that it increases the length of the prompt, which may exceed the token limit of certain models.
For few-shot prompting, the best-performing model is Claude 3.5 Sonnet. It outperforms all other models in terms of average performance across both categories. It achieves over 98% in precision, recall, and F1 score for the valid category, and over 86% for invalid category. However, applying the self-consistency method to Claude 3.5 Sonnet does not significantly improve its overall performance, as the F1 score improvement (in percentage) in both categories remains within 3%.
Figure 5. F1 Score Comparison for Euclidean and Cosine Similarity Across Different Values of K.
[…]
For RAG, we first compare the performance of TVR with different Ks (i.e., the number of examples provided in the prompt). Figure 5 illustrates the F1 score curves for both categories as K varies, using Euclidean and cosine similarity measures. The results show that initially, as K increases, TVR’s performance improves in both categories, with a particularly significant improvement in the invalid category. When K reaches 3, the F1 score peaks for both categories, with the cosine similarity yielding a slightly higher F1 score than Euclidean distance. However, as K continues to increase, TVR’s performance slightly decreases. This may result from an excessive number of examples, including dissimilar ones, that are potentially confusing the model and thus lowering accuracy. Therefore, the optimal configuration of RAG-based TVR is using cosine similarity with K=3.
In conclusion, TVR using Claude 3.5 Sonnet with RAG achieved the best performance in requirements traceability validation, with the highest accuracy of 98.87% across all configurations. It also demonstrated high precision, recall, and F1 scores, reaching over 99% for valid traceability links and over 93% for invalid ones, highlighting the effectiveness in retrieving similar examples and including them in the prompt, thus helping the LLMs to better understand the requirement pair to be verified.
Answering RQ2: Claude 3.5 Sonnet with RAG achieved the best performance in requirements traceability validation when using cosine similarity with K set to 3. Its high accuracy makes it a viable solution for traceability validation in practice.
5.3. Robustness to Unseen Requirements Variations (RQ3)
Approach. This RQ evaluates the robustness of TVR to unseen variations of stakeholder requirements, a crucial aspect in practice. To ensure a comprehensive evaluation, we employ cross-validation by categorizing all requirement pairs based on their variations and evaluating across all four categories in our dataset. Specifically, for each requirement pair to be validated, our retriever component first identifies its variation category and then retrieves the most similar examples from the remaining three variation categories, thus emulating situations where new variations are encountered.
Results. Table 2 presents the TVR robustness evaluation results. Accuracy results suggest TVR’s strong generalization ability across variation categories for valid pairs, with an F1 score of 98.43%. However, TVR struggles with invalid pairs, exhibiting a lower recall of 74.88%.
Acc(%) Valid Invalid
Pre(%) Recall(%) F(%) Pre(%) Recall(%) F(%)
All 97.13 97.40 99.48 98.43 93.83 74.88 83.29
V1 93.96 94.37 98.75 96.51 90.79 67.65 77.53
V2 98.78 98.81 99.89 99.35 98.28 83.82 90.48
V3 98.44 98.59 99.76 99.18 95.24 76.92 85.11
V4 92.59 95.00 95.00 95.00 85.71 85.71 85.71
Table 2. Robustness Evaluation Results.
Among variation categories, V4 shows the lowest accuracy (92.59%), with F1 scores of 95.00% for valid pairs and 85.71% for invalid pairs. V1 follows with an accuracy of 93.96%, showing a significant drop in recall for invalid pairs (67.65%), suggesting frequent misclassifications of invalid pairs as valid. In contrast, results for V2 and V3 are similar, with high accuracy rates of 98.78% and 98.44%, respectively. For V2, TVR demonstrated strong adaptability, excelling in both valid (F1 score: 99.35%) and invalid pairs (F1 score: 90.48%), whereas TVR achieved a lower recall for the invalid pairs (76.92%) of V3.
These results may be attributed to two factors: (1) Variability in category differences affects the quality of retrieved examples in guiding TVR. Specifically, V1 belongs to the “Lost Communication” type of DTCs, whereas the remaining three variations fall under the “Implausible Data” category. This discrepancy likely contributed to TVR’s poorer prediction performance during cross-validation for V1. (2) Certain variations exhibit greater complexity, making them more challenging for the LLM to understand and predict accurately. Although V2, V3, and V4 all belong to the “Implausible Data” category, our analysis revealed that V4 is more complex than other variations, containing more messages and conditions to be verified. Moreover, the template for V4 is not strictly fixed, leading to slightly lower accuracy during cross-validation. Furthermore, V4 has only 27 cases, resulting in a small denominator in the evaluation metrics, and thus lower reliability in results.
Answering RQ3: TVR achieved an overall accuracy of 97.13% on robustness evaluation, indicating its strong generalization ability across unseen variation categories, with more challenges for certain variation categories that are more distinct and invalid pairs.
5.4. Performance on Requirements Traceability Recovery (RQ4)
Approach. As explained in Section 3.3, we employ a three-step pre-filtering process to reduce the number of requirement pairs we consider could have a missing link. For the remaining pairs, TVR predicts whether there is a valid traceability link between them. We then calculated the Correctness of pairs predicted as ‘Yes’.
Results. The three stages of pre-filtering yielded, in turn, 30,598, 14,494, and 1,919 requirement pairs. This demonstrates the effectiveness of our three-step pre-filtering, which significantly reduces infeasible pairs. Then 502 pairs out of 1,919 pairs are predicted to have valid traceability links. After manually verifying all 502 pairs, we obtained a Correctness of 85.50%, suggesting that most retrieved links are correct, with a limited number of false positives.
We conducted an error analysis and found that all 73 prediction errors stemmed from the same issue. In the stakeholder requirement, TVR first verifies whether the trigger condition is set to “RUN”, followed by checking the input message regarding setting or clearing the DTC. Since a trigger condition is also a message, if the corresponding system requirement includes the same message, the model may incorrectly interpret that there is a traceability link, whereas it should actually be checking for the input message. To enhance retrieval accuracy, future approaches can further improve by identifying and excluding the trigger condition while retaining only the content related to the input message.
Answering RQ4: After effectively reducing the number of requirements pairs being considered with our three-step filtering approach, TVR achieves an 85.50% correctness in recovering missing links, thus making it applicable with a manageable number of false positives. Based on an error analysis, we conclude that additional preprocessing of messages and signals may further improve accuracy.
6. Discussion
6.1. Industry Feedback on TVR
Industry feedback on TVR underscores its significant advantages. Engineers were impressed by its high accuracy and reliability, thus recognizing its potential for ensuring traceability consistency in practice. TVR’s high level of automation can significantly reduce manual effort, leading to substantial time and cost savings.
Furthermore, we demonstrated TVR’s interpretable output to system engineers by removing the instruction “and only respond with either ‘Yes’ or ‘No’” from our prompt. In its response, TVR first identifies the key message and signal in the stakeholder requirement. Then it searches the relevant message or signal in the linked system requirement and checks whether the system requirement addresses the message or signal in the stakeholder requirement by taking consistent action for it. For an invalid pair, TVR provides an example explanation as follows:
“Step 1: Identify the key message/signal in the stakeholder requirement. The key message is MESSAGE_1. Step 2: Search for this message in the system requirement. After carefully examining the system requirement, I don’t see any explicit mention of MESSAGE_1 or any checks related to it. Step 3: Determine if the message is covered. Since there is no explicit mention or check for MESSAGE_1, we can conclude that this specific message is not covered. Step 4: Conclusion. The response is: No.”
This interpretable output enables system engineers to quickly understand and validate traceability decisions, thereby enhancing TVR’s applicability.
Looking ahead, industry professionals identified two key points for improvement: Further minimizing false positives to further refine precision and enhancing scalability to better adapt to large-scale industrial datasets with diverse requirements. These insights will inform future refinements, ensuring the approach remains both effective and practical for real-world adoption.
6.2. Practical Application of TVR
TVR is designed to automate and enhance the accuracy of traceability link validation and recovery for DTC requirements in automotive systems. To apply TVR, practitioners only need to prepare three key components: (1) The input pairs to be verified – stakeholder requirements and system requirements. (2) The retrieval database – a collection of requirement pairs with labels (valid or invalid) previously verified by system engineers. (3) The prompt specifying the task description and instructions. Once these inputs are provided, TVR generates a prediction result with an explanation as instructed by the prompt, improving transparency and interpretability.
TVR is particularly well-suited for iterative software development in industrial settings. As the retrieval database expands with more human-confirmed data, the retriever can extract more relevant examples, enhancing the LLM’s ability to make more accurate and reliable assessments over time.
When applying TVR to different datasets (beyond DTC requirements) or various software artifacts, the prompt instructions must be adjusted to ensure the task description aligns closely with the specific problem. Additionally, selecting the appropriate number of examples and the similarity measure requires empirical experimentation with the specific data at hand to determine the optimal configuration. A priori, the basic principles of TVR can also be adapted to other domains and types of requirements.
In summary, TVR demonstrates excellent performance on industrial data, offering high accuracy, strong robustness, ease of use, and broad applicability.
7. Threats to Validity
In this study, we identify the following threats to validity.
External Validity. TVR is specifically designed for DTC requirements, which are critical in the automotive domain, using a dataset from an industry partner. Since TVR relies on a retrieval-based approach to obtain prompt examples, the accuracy of TVR in practical applications may be affected by the quality of the retrieval database. However, the general principles of TVR should be widely applicable to many types of requirements in domains where traceability between high-level stakeholder requirements and system requirements is important.
Internal Validity. The dataset was annotated by co-authors under the guidance of system engineers. However, manual annotation is inherently prone to errors, such as misinterpretation of requirements or inconsistencies in labeling. These annotation errors could affect the evaluation of TVR’s performance. Another internal threat arises from the model’s robustness evaluation, which is affected by the degree of variation among different categories. To mitigate this threat, we applied the concept of cross-validation to evaluate the model’s robustness across all four categories.
8. Related Work
Traceability, defined as “the ability to describe and follow the life of an artifact developed during the software lifecycle in both forward and backward directions”  (Lucia et al., 2007), plays a crucial role in requirements engineering and software engineering in general. To date, the most extensively studied challenges are traceability link recovery and traceability maintenance (Guo et al., 2024). Researchers have proposed various approaches to support software traceability between different artifacts, including requirements-to-code (Gao et al., 2022; Hey et al., 2021; Panichella et al., 2013; Kuang et al., 2015), document-to-code (Antoniol et al., 2002; Marcus and Maletic, 2003; Keim et al., 2024), and document-to-model (Keim et al., 2023; Cleland-Huang et al., 2005).
Early automated traceability techniques relied on classical NLP methods, primarily leveraging information retrieval techniques (Guo et al., 2024). These methods establish potential traceability links by computing textual similarity between artifacts using models such as the vector space model (Antoniol et al., 2002; Mahmoud, 2015), latent semantic indexing (LSI) (Marcus and Maletic, 2003), latent Dirichlet allocation (LDA) (Asuncion et al., 2010), and hybrid approaches (Gethers et al., 2011; Moran et al., 2020). However, these traditional models struggle to capture deep semantics. Recent approaches have addressed this limitation by incorporating word embeddings and deep neural networks (Guo et al., 2017; Wang et al., 2018; Zhao et al., 2017; Hey et al., 2021). Additionally, some studies have explored active learning (Mills et al., 2019) and self-attention mechanisms (Zhang et al., 2021) for improving traceability link recovery.
With the advent of LLMs such as GPT (OpenAI, 2023a, b), Llama (https://www.llama.com/, 2023), and Claude (https://claude.ai/, 2023), new opportunities have emerged for automating traceability while addressing the limitations of previous methods. Hey et al. (Hey et al., 2021) proposed the FTLR approach, which integrates word embedding similarity and Word Mover’s Distance (Kusner et al., 2015) to generate requirements-to-code trace links. Hey et al. (Hey et al., 2024) further enhances accuracy by incorporating a NoRBERT classifier to filter irrelevant requirement segments. Rodriguez et al. (Rodriguez et al., 2023) investigated the use of Claude, to directly perform traceability link recovery through natural language prompting. Their study demonstrated that generative models can not only recover trace links but also provide explanations for identified links. Fuchss et al. (Fuchß et al., 2025) recently introduced LiSSA, a RAG-enhanced LLM approach for traceability link recovery across requirements-to-code, documentation-to-code, and architecture documentation-to-models. LiSSA first retrieves relevant elements based on similarity values between embeddings, then generates LLM-based prompts to assess traceability links.
In contrast to prior studies, our research focuses on validating traceability links as well as recovering missing links, addressing specific industry needs in the automotive sector: 1) Industry practitioners require a mechanism to verify the correctness of traceability links between stakeholder requirements and system requirements established by system engineers. 2) Both stakeholder and system requirements exhibit variations in the way they are expressed, with the possibility of encountering unseen variations in the future. 3) Both stakeholder and system requirements follow loose templates, with small differences in the way message and signal values are expressed, thus rendering information retrieval-based approaches based on term frequency and word overlapping ineffective. To address these challenges, we frame the validation task as a binary classification problem, where LLMs determine the validity of existing traceability links. LLMs are well-suited for this task due to their ability to understand and capture the subtle differences in messages and signals, in the presence of values among variations in requirements templates while simultaneously supporting explainability for traceability link validation.
9. Conclusion
In this paper, we propose TVR, an approach leveraging RAG based on LLMs to verify the validity of traceability between high-level stakeholder requirements and system requirements in automotive systems. TVR achieves 98.87% accuracy in detecting whether a traceability link is valid. Furthermore, experimental results demonstrate the robustness of TVR in effectively handling unseen variations in requirements templates, retaining a 97.13% accuracy. Additionally, TVR can identify missing links between requirements with 85.50% correctness. These findings indicate that TVR is effective not only for traceability validation but also for recovering missing links. TVR can thus be applied to automotive systems, helping the industry to save both time and cost. In the future, the basic principles of TVR can be adapted to other types of requirements and domains where such traceability is important.
Acknowledgements.
This work was supported by a research grant from Aptiv, as well as by the Discovery Grant and Canada Research Chair programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Science Foundation Ireland grant 13/RC/2094-2.